OpenStudy (anonymous):

You roll two dice. What is the probability that the sum of the dice is greater than 5?

OpenStudy (anonymous):

26/36 = 13/18?

hartnn (hartnn):

(1,5),(1,6)--->2 (2,4),(2,5),(2,6)--->3 and so on... total = 2+3+4+5+6+6=26 probability = 26/36 = 13/18 yes , you are correct :)